# Bridging Voting and Deliberation with Algorithms: Field Insights from vTaiwan and Kultur Komitee

### v1

@@ 5.3. The ReadTheRoom Deliberation Method
5.3.1. Voting shows how deliberation reduces polarisation and builds consensus
Using the decision tree in Figure 10, participants started deliberating the top five statements that divided the online opinion, gathered using Polis beforehand. Figure 11 illustrates the impact of deliberation on participant opinions across the five statements, providing insights into opinion shifts using four key metrics. The percentage of voters changed shows significant variability, with 53.1% of participants changing their stance on Statement 2 (AI companies should disclose algorithms and data), the highest among the statements. In contrast, only 30.0% of participants changed their opinion on Statement 3 (AI’s industrial impact should be solved by the market), indicating more entrenched positions. The Polarisation Index highlights how deliberation reduced opinion dispersion, particularly for Statement 2 (0.75), whereas Statement 3 retained high polarisation (0.95), reflecting persistent disagreement. The Consensus Index reveals increased alignment for most statements; for example, Statement 5 (AI developers must be accountable for harm caused by AI content) improved from 0.44 to 0.47, while Statement 4 (Governments should prevent AI reinforcing stereotypes) rose from 0.41 to 0.43. Finally, Mean Opinion Changes capture shifts in group sentiment, such as Statement 4 shifting significantly from -0.17 to 0.69, demonstrating a strong movement toward agreement. Conversely, Statement 2 shifted slightly toward disagreement (-0.03 to -0.34). Together, these metrics reveal that deliberation effectively fosters opinion shifts, reduces polarisation, and builds consensus, though its impact varies depending on the statement and participants’ initial positions.
5.3.2. Participants were able to voice their diverse opinions with ReadTheRoom method
In the survey (Figure 17 in the appendix) conducted at the end of the deliberation with 5-point Likert scale (-2 to 2), the ReadTheRoom Deliberation method demonstrates its strength in fostering mutual learning among participants. The highest mean (1.60) for “The discussion today helped me learn new perspectives” highlights the process’s effectiveness in broadening participants’ understanding. The positive responses suggest that the deliberation created an environment conducive to knowledge exchange and mutual understanding. Additionally, the significant agreement on “real-time voting” (a core aspect of the ReadTheRoom method) allowing diverse views to be expressed (1.57) and the constructive consensus-building process (1.13) further illustrates the method’s ability to encourage participants to reflect on diverse perspectives and adapt their views. By visualising opinion changes transparently, the deliberation fosters a shared learning experience where participants collectively engage with divisive issues and work toward consensus. Most participants also see the point of having a more data-supported deliberation outcome, as they agree that the combination of voting and deliberation makes the outcomes more convincing (1.47).
Feedback from participants also shows that the method was able to prompt meaningful dialogue. P05 remarked, “There were many deep experiences and opinions from various fields!”. P08 noted, “Great, the data generated today is very valuable!”. P10 expressed, “I learned a lot about different positions and their concerns, and I gained a lot!”. P13 shared, “My understanding of related issues became much clearer”. Some feedback recommended that vTaiwan use even more digital tools and demand more from citizen input. P21 suggested, “I recommend asking participants to also document changes in their attitudes simultaneously”. While P39 proposed, “More digital tools can be used to continue the dialogue online for deeper discussions in the future”.
@@ 6.1. Practical Impact
This study examined three digital methods — Radial Clustering for Preference Based Subgroups, Human-in-the-loop MES, and ReadTheRoom Deliberation — implemented in two distinct participatory processes. While conducted on a relatively modest scale, these method offers innovative enhancements with broader implications.
First, Radial clustering facilitated the creation of small, evenly sized breakout groups during the deliberation phase of KK24. Differences in outcomes between deliberation rounds underscore the impact of varying group compositions on decision-making dynamics. This method aligns with the emphasis on bridging enclaves with the wider deliberative process (Sunstein, 2017).
Crucially, this approach demonstrates how Enclave Deliberation can be achieved without stereotyping or profiling citizens, relying solely on real preferences derived from online voting data. This aligns with Abdullah et al. (2016)’s emphasis that enclaves should “represent marginalised perspectives or social locations rather than essentialised identities”. Such an approach avoids reductive assumptions about individuals’ views based on their demographic characteristics, instead empowering disadvantaged groups to contribute autonomously and effectively to broader deliberations. Future research could examine the impact of this clustering approach on the quality and outcomes of deliberation in larger-scale settings.
@@ 6.2. Limitations
Throughout the application and research, we identified limitations affecting both the methods and their impact. One overarching limitation of this study is its reliance on case studies with small sample sizes. While these offer valuable insights, they limit the generalisability of findings and may not fully capture the complexities of larger-scale applications. However, we chose this approach because these methods are highly relevant for real-world situations, where small group settings and practical constraints often shape deliberative processes. Future research should validate these methods in more diverse and extensive settings to ensure broader applicability.
For group composition with Radial clustering, the quantitative differences in agreement between the two conditions (homogeneous and heterogeneous) could have been influenced by confounding factors such as gender and age. Group demographics may have impacted dynamics beyond the intended effects of voting-based clustering. Additionally, the algorithm’s use of a low-dimensional space (e.g., PCA) prioritised visual clarity over precision, potentially losing critical information. While the method ensures equal group sizes, future work should refine the algorithm to balance transparency with and clustering accuracy.
Human-in-the-loop MES allowed flexibility in adjusting budget ratios and modifying individual project budgets, but some challenges were identified. The KK24 organisers set a 50:50 default budget ratio, potentially creating a default bias. Although participants considered alternatives, they ultimately retained the default. Future research should explore whether removing preset defaults changes decision-making outcomes. Additionally, while participants could reduce some project budgets, this flexibility might compromise the proportionality MES aims to ensure, potentially leading to a sub-optimal allocation of resources. Pre-adjusting project budgets could enhance fairness, but it may also place an excessive burden on participants.
In the ReadTheRoom Deliberation, participants voted on the same statements twice: before and after the discussion. Some participants raised concerns about the definitions and wording of these statements, suggesting that updating them in real-time to reflect ongoing conversations could improve the process. A potential future direction could involve using AI-driven summarisation of deliberations based on real-time transcripts to dynamically update the statements. This approach could make responses more relevant and insightful. While the interactive design effectively fosters discussion, a notable limitation of real-time visualisation is its potential to influence individual voting, creating a bias toward prevailing opinions. To mitigate this, incorporating some level of secret voting could be essential, particularly in high-stakes political scenarios.
@@ 6.3. Concluding Remarks
Deliberation and voting are not opposing endpoints but rather complementary points on a spectrum of democratic innovation (Weyl et al., 2024). By combining computational methods like Radial clustering, Human-in-the-loop MES, and ReadTheRoom, this study shows how algorithms can bridge the gap between deliberation and voting, addressing the trade-offs between depth and breadth in democratic processes. The advancements presented here point toward a future where richer collaboration and broader participation are not mutually exclusive; algorithms and thoughtful design can help democratic innovations overcome traditional constraints for fairer, more impactful decision-making.
@@ Acknowledgements.
We thank the entire KK24 committee and vTaiwan community, as well as all the citizens who participated, for their invaluable contributions to this research and to participatory democracy. Special thanks to Noemi Scheurer and Mia Odermatt, the main organisers of Kultur Komitee Winterthur, for their openness to experimenting with innovative methods. We also thank Jia-Wei Cui, the current organiser and moderator of vTaiwan, for his commitment, and Yi-Ting Lien for co-hosting these deliberation workshops and connecting vTaiwan with TWNIC. Additionally, FB gratefully acknowledges the financial support of the Swiss National Science Foundation (SNSF) under grant ID CRSII5-205975.References
@@
- Faliszewski et al. (2023) Piotr Faliszewski, Jarosław Flis, Dominik Peters, Grzegorz Pierczyński, Piotr Skowron, Dariusz Stolicki, Stanisław Szufa, and Nimrod Talmon. 2023. Participatory budgeting: data, tools, and analysis. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence (Macao, P.R.China) (IJCAI ’23). Article 297, 8 pages. https://doi.org/10.24963/ijcai.2023/297
@@
- Gastil and Broghammer (2021) John Gastil and Michael Broghammer. 2021. Linking Theories of Motivation, Game Mechanics, and Public Deliberation to Design an Online System for Participatory Budgeting. Political Studies 69, 1 (2021), 7–25. https://doi.org/10.1177/0032321719890815
@@
- Witting et al. (2025) Irene Witting, Charlotte Wagenaar, and Frank Hendriks. 2025. Improving referendums with deliberative democracy: A systematic literature review. International Political Science Review 46, 1 (2025), 40–56. https://doi.org/10.1177/01925121231210048
- Yang et al. (2024) Joshua C. Yang, Carina I. Hausladen, Dominik Peters, Evangelos Pournaras, Regula Hnggli Fricker, and Dirk Helbing. 2024. Designing Digital Voting Systems for Citizens: Achieving Fairness and Legitimacy in Participatory Budgeting. Digit. Gov.: Res. Pract. 5, 3, Article 26 (Sept. 2024), 30 pages. https://doi.org/10.1145/3665332
@@
7. Appendices
Appendix A Radial Clustering for Preference Based Subgroups
| Metric | Homogeneous | Heterogeneous |
| Total Budget | 186,900 | 189,900 |
| Number of Projects | 14 | 10 |
| Mean Cost | 13,350.00 | 18,990.00 |
| Median Cost | 11,450.0 | 13,000.0 |
| Cost Std. Dev. | 8,581.71 | 12,657.05 |
| Mann-Whitney U | 49.0 (p-value: 0.22715) | |

### v2

@@ 5.3. The ReadTheRoom Deliberation Method
5.3.1. Deliberation reduced polarisation and built consensus
Using the decision tree (Figure 12) in the appendix, participants deliberated on the five most divisive statements from the Polis online survey. Our analysis of before-and-after voting (Table 2) reveals consistent patterns across all five statements deliberated. Between 30.0% and 53.1% of participants changed their stance following deliberation, with the highest shift observed for Statement 2 on AI company transparency (53.1%) and the lowest for Statement 3 on market-based solutions (30.0%). The Bimodality Coefficient (BC) decreased universally after deliberation, with particularly notable reductions for Statement 2 (from 0.561 to 0.344, a 38.7% decrease) and Statement 5 on developer accountability (from 0.526 to 0.443, a 15.8% decrease). Importantly, three statements began with BC values near or above the polarisation threshold of 0.555, but all fell below this threshold after deliberation, indicating a clear shift from divided opinion camps toward more cohesive viewpoints. This reduction in polarisation was complemented by increases in the Consensus Index across all statements, with Statement 2 showing the largest improvement (from 0.43 to 0.50). These consistent gains in consensus suggest that deliberation effectively brought participants toward shared understanding, even on initially divisive topics. Mean opinion values showed varied directional changes. Statement 4 on government prevention of AI bias demonstrated the most dramatic and only statistically significant shift, moving from slight disagreement (-0.17) to clear agreement (0.69). The limited statistical significance in our dataset (with only one change reaching p ¡ 0.05) likely stems from our modest sample size rather than the absence of an effect. The consistency of changes across multiple metrics—decreasing BC values, increasing Consensus Indices, and substantial voter position changes—provides evidence that the ReadTheRoom method facilitates productive opinion convergence whilst preserving necessary distinctions in perspective.
5.3.2. Participants valued the learning experience and transparency of the process
Post-deliberation survey responses (Figure 15) highlight the ReadTheRoom method’s effectiveness. On a 5-point Likert scale (-2 to 2), participants gave the highest rating (1.60) to “The discussion today helped me learn new perspectives,” confirming the method created a genuine learning environment. Participants strongly agreed (1.57) that real-time voting allowed diverse views to be expressed effectively. They also valued the consensus-building process (1.13) and found the combination of voting with deliberation convincing (1.47). Qualitative feedback reinforces these positive assessments. P05 noted, “There were many deep experiences and opinions from various fields!”, whilst P10 shared, “I learned a lot about different positions and their concerns, and I gained a lot!” P13 reported a clearer understanding of the issues, and several participants suggested expanding the digital tools further. P39 proposed using “more digital tools to continue the dialogue online for deeper discussions”. These findings demonstrate how the ReadTheRoom method transforms potentially divisive discussions into collaborative learning experiences where participants engage meaningfully with different perspectives. By making opinion shifts visible in real time, the process fosters mutual understanding whilst providing policymakers with data-supported insights into public sentiment.
@@ 6.1. Practical Impact
This study examined three digital methods—PCD, Human-in-the-loop MES, and ReadTheRoom Deliberation—implemented in two participatory processes. While conducted on a modest scale, these methods offer innovative enhancements with broader implications.
First, the PCD framework offers a fresh approach to structuring deliberative participation. By using voting data to form discussion groups, it creates a practical bridge between individual preferences and collective decision-making. Our findings from KK24 show this method impacts both the process and outcomes in meaningful ways. Participants in homogeneous groups reported finding decisions easier to make, with outcomes better reflecting their preferences. This suggests that bringing similar voters together creates a supportive environment where niche interests can be fully expressed before broader engagement. Interestingly, our analysis showed that homogeneous groups tended to fund more numerous, smaller projects, while heterogeneous groups favoured fewer, costlier initiatives. This pattern suggests homogeneous deliberation may naturally support diverse smaller projects serving specific interests, whereas heterogeneous groups converge on larger projects with broader appeal as participants seek common ground across differing priorities. This approach aligns with Sunstein’s (Sunstein, 2017) concept of enclave deliberation, while the later heterogeneous phase helps prevent potential preference polarisation. A significant strength of PCD is how it implements enclave deliberation without using demographic assumptions. As Abdullah et al. (Abdullah et al., 2016) emphasise, democratic enclaves should “represent marginalised perspectives or social locations rather than essentialised identities”. While we used Radial Clustering for its visual clarity and balanced groups, the PCD framework remains flexible. Different clustering methods could be selected based on specific deliberative goals and contexts. The visual simplicity of our approach had practical benefits, making the grouping logic easily understood by facilitators and participants without technical backgrounds.
@@ 6.2. Limitations
Our research faced several methodological and contextual limitations. While our case studies offered valuable real-world insights, the modest sample sizes limit statistical power and generalisability. These participant numbers, though typical for deliberative mini-publics, may not capture the complexity of dynamics present in larger democratic innovations or different cultural contexts.
For the PCD framework, our findings suggest that clustering interacts with demographic factors in ways that affect deliberative outcomes. Social dynamics like confidence levels, speaking styles, and status hierarchies likely influence how different groups benefit from various deliberation structures. This makes it difficult to isolate preference-based effects from intersecting social factors. Additionally, as the KK24 outcomes represented aggregated results combining points from all groups across both deliberation rounds, it was more difficult for the participants to precisely identify the direct influence of different deliberation structures on final outcomes. Our implementation of Radial Clustering prioritised visual interpretability and practical usability over mathematical optimisation. While achieving reasonable silhouette scores, this approach potentially sacrificed precision in capturing the full dimensionality of preference spaces. The two-dimensional PCA projection, though visually intuitive, inevitably flattens the complexity of participant preferences.
With Human-in-the-loop MES, participants reviewed and modified project budgets after the MES calculation. These post-selection modifications potentially compromised MES’s mathematical proportionality guarantees, creating tension between algorithmic fairness and participant agency. While such adjustments often incorporate valuable contextual knowledge, they distort the carefully calculated proportional representation that makes MES theoretically fair. Future research should develop extensions to MES that formally account for these post-selection budget modifications while preserving core fairness properties, treating such adjustments as preference updates within the algorithm itself.
As for the ReadTheRoom method, participants noted that fixed statement wording couldn’t evolve alongside the discussion, potentially constraining the natural progression of ideas. The visible spectrum-based voting display might have introduced social conformity effects, with participants potentially adjusting their positions to align with perceived group norms rather than expressing independent judgments. Additionally, our evaluation captured only immediate opinion shifts during a single deliberative session, without measuring the stability of these changes over time. The observed reductions in polarisation might represent temporary accommodations rather than durable opinion changes.
Also, our evaluation approaches across all three methods prioritised practical measurement in field settings over controlled experimental design. The context-specific metrics we developed lack standardisation that would facilitate direct comparison with other deliberative innovations. Future work should develop more standardised evaluation frameworks that balance theoretical rigour with practical applicability in democratic contexts.
@@ 6.3. Concluding Remarks
Deliberation and voting are not opposing endpoints but complementary facets of democratic innovation (Weyl et al., 2024). Our research demonstrates how computational methods can bridge deliberative depth with participatory breadth by transforming voting data from simple preference aggregation into a resource for structuring meaningful group interactions. The PCD framework, Human-in-the-loop MES, and ReadTheRoom approach collectively show how thoughtfully designed algorithms can help democratic processes overcome traditional constraints, creating decision-making systems that are simultaneously more representative, reflective, and effective as communities face increasingly complex collective challenges.
@@ Acknowledgements.
We thank the KK24 committee and vTaiwan community, as well as all participants, for their invaluable contributions to this research on participatory democracy. Special thanks to Noemi Scheurer and Mia Odermatt, the main organisers of Kultur Komitee Winterthur, for their openness to experimenting with innovative methods. We also thank Jia-Wei Cui, the current organiser and moderator of vTaiwan, for his commitment, and Yi-Ting Lien for co-hosting these deliberation workshops and connecting vTaiwan with TWNIC. Additionally, FB gratefully acknowledges the financial support of the Swiss National Science Foundation (SNSF) under grant ID CRSII5-205975.References
@@
- Faliszewski et al. (2023) Piotr Faliszewski, Jarosław Flis, Dominik Peters, Grzegorz Pierczyński, Piotr Skowron, Dariusz Stolicki, Stanisław Szufa, and Nimrod Talmon. 2023. Participatory budgeting: data, tools, and analysis. In Proceedings of the Thirty-Second International Joint Conference on Artificial Intelligence (Macao, P.R.China) (IJCAI ’23). Article 297, 8 pages.
@@
- Gastil and Broghammer (2021) John Gastil and Michael Broghammer. 2021. Linking Theories of Motivation, Game Mechanics, and Public Deliberation to Design an Online System for Participatory Budgeting. Political Studies 69, 1 (2021), 7–25.
@@
- Knapp (2007) Thomas R Knapp. 2007. Bimodality revisited. Journal of Modern Applied Statistical Methods 6, 1 (2007), 3.
@@
- Witting et al. (2025) Irene Witting, Charlotte Wagenaar, and Frank Hendriks. 2025. Improving referendums with deliberative democracy: A systematic literature review. International Political Science Review 46, 1 (2025), 40–56.
- Yang et al. (2024a) Joshua C. Yang, Carina I. Hausladen, Dominik Peters, Evangelos Pournaras, Regula Hänggli Fricker, and Dirk Helbing. 2024a. Designing Digital Voting Systems for Citizens: Achieving Fairness and Legitimacy in Participatory Budgeting.
- Yang et al. (2024b) Joshua C. Yang, Carina I. Hausladen, Dominik Peters, Evangelos Pournaras, Regula Hnggli Fricker, and Dirk Helbing. 2024b. Designing Digital Voting Systems for Citizens: Achieving Fairness and Legitimacy in Participatory Budgeting. Digit. Gov.: Res. Pract. 5, 3, Article 26 (Sept. 2024), 30 pages. https://doi.org/10.1145/3665332
@@
Appendix A Preference-based Clustering for Deliberation
| Algorithm | SS | BG | OC |
| Radial Clustering | 0.238 | Yes | No |
| Low-Dim Balanced K-Means | 0.220 | Yes | No |
| High-Dim Balanced K-Means | -0.051 | Yes | Yes |
| One-Dim PCA Clustering | 0.030 | Yes | No |
| Low-Dim K-Means | 0.429 | No | No |
| High-Dim K-Means | -0.175 | No | Yes |